if the area of a kite is 150mm² and the longer diagonal has length thrice the shorter diagonal. What are the length of its diagonals?
1
Expert's answer
2010-07-28T04:21:31-0400
Area of kite = (1/2)*D1*D2. Let D1=x, D2=3x, so 3x^2=150*2=300. x^2=100, x=10. Thus the diagonals are D1=10, D2=30.